OpenStudy (anonymous):

Help! 5x/x+1=4x/x-1+10/x^2-1

OpenStudy (anonymous):

\[\frac{5x}{x+1}=\frac{4x}{x-1}+\frac{10}{x^2-1}\] \[\frac{5x}{x+1}=\frac{4x}{x-1}+\frac{10}{(x-1)(x+1)}\] \[(x-1)(x+1)(\frac{5x}{x+1})=(x-1)(x+1)(\frac{4x}{x-1})+(x-1)(x+1)(\frac{10}{(x-1)(x+1)})\] \[\frac{5x(x-1)}{(x-1)(x+1)}=\frac{4x(x+1)}{(x-1)(x+1)}+\frac{10}{(x-1)(x+1)}\] \[5x(x-1)=4x(x+1)+10\] \[5x^2-5x=4x^2+4x+10\] \[x^2-9x-10=0\] \[(x-10)(x+1)=0\] \[x=10, -1\] Because -1 is not in the domain of this rational equation, x=10 is your only solution

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw